# Multi-view Symbolic Regression

### v2

@@ 5.2. Finance dataset
Financial markets exhibit complex emergent behaviours. We define the return as the difference between the price of an asset at time $t$ and time $t+1$ is $r_{t}=p_{t+1}-p_{t}$. At first approximation, we can describe the distribution of returns by a Brownian motion (Bachelier, 1900). The field of econophysics, which applies tools from statistical physics to study these stochastic processes, seeks to identify probability distribution functions that go beyond such Gaussian approximation. Those initial models led to the development of the renowned Black-Scholes equation (Black and Scholes, 1973), but neglected the significance of rare and extreme events due to their lack of fat tails. Indeed, the distribution of these specific events can be fitted with power-laws (Mantegna and Stanley, 1995).
@@
Modern models, as presented in (Mantegna and Stanley, 1999), focus on Lévy processes. These include, for example, the Gaussian distribution as well as the Cauchy distribution, the second being an already improved solution with its fat tails. Such models were proposed after studying the statistical properties that most data exhibit, such as the famous $S\&P500$ dataset, and then finding distributions that possess such properties. However, our Multi-view approach enables us to identify a common distribution for all assets by considering each of them individually. In this Section, we show that MvSR rediscover some of the presented distributions and propose new models that better fit the real return distributions.
@@
We analyse time series data consisting of multiple prices generated by various assets. Each asset’s price corresponds to the value of a share of a company on financial markets. We utilise a publicly available Kaggle dataset  [^4] https://www.kaggle.com/datasets/iveeaten3223times/massive-yahoo-finance-dataset containing data from $491$ companies. The first 10 companies are used as views for MvSR, while the remaining companies are only used for testing purposes. Values of assets are taken each day at open market time, over a period of 5 years starting from January $1^{st}$, $2018$. For each asset, we study the distribution of its returns with a sampling of $100$ bins of equal width. We normalize the data as described in Section 3.1. This type of data exhibits common statistical properties. A positive mean that corresponds to overall economic growth. A leptokurtic profile that indicates more extreme events than a normal distribution would produce, and a negative skewness due to an asymmetric shape that informs rare events are more likely to be crises than economic booms. The first and last properties are only observed for datasets covering long time periods (months or years) such as ours.
@@
Under those conditions, we explore multiple seeds, tree lengths (ranging from size 8 to 20) and operators (using various combinations of $exp$, $abs$ and $power$) to obtain multiple possible parametric solutions with MvSR. In Table 3 we present 6 parametric models generated by MvSR. We measure the performance of all the functions by fitting them to each asset and display the median MSE value as a comparative metric in the last two columns of Table 3.
@@
Figure 4. Normalized (Section 3.1) distribution of returns for 2 assets, fitted by the Cauchy model and the best MvSR solution (Power-Laplace).
Figure 4 shows the distributions of returns of some assets alongside the fit given by the best distributions, new and already found in the literature, that we obtain. We call the latter a Power-Laplace distribution  [^5] That we name after it’s functional composition $h(x)=g(f(x))$ with the Laplace distribution $g(x)=e^{-a\times|x|}$ and a power function $f(x)=x^{b}$.. It performs well because the power is always a root  [^6] We always find $10<b<$ 1., which accentuates the fat tail effect of the distribution, allowing a better fit in the tails while maintaining a good peaked fit in the centre.
@@
Finally, we see that distributions found in the literature are relevant for the $S\&P500$, as they all perform much better than the Gaussian. However, apart from the Gaussian, their scores do not differ much. This shows that they are almost all equivalent fits for this aggregated index. On the other hand, their median scores for a variety of assets are much more disparate. In this second case, our algorithm produced much more efficient distributions ; distributions that outperformed all the others. This indicates that our multi-view approach is an efficient strategy for finding a unique distribution that gives good fits for as many assets as possible. We illustrate this effect on Figure 4. The right panel shows how Cauchy and Power-Laplace both fit well the $S\&P500$ index. The left panel displays a special case selected for its high score difference between both models. This shows how a model like Cauchy fails to fit some distributions compared to the best model found by MvSR, which is more versatile.
We showed that applying MvSR to a set of assets all at once, rather than looking at them separately or using an aggregated index, is an efficient approach to gaining new insights into markets behaviour and improving characterisation of the stochastic process that governs them. The results presented here could be fine tuned by adding more functions to the MvSR. For example the Gamma function could allow us to generate more distributions such as the Gamma variance process (Madan et al., 1998) and Student’s t-distribution (Bouchaud and Potters, 2000), which are the general forms for the distributions we already found using MvSR. Finally, further investigation of the distributions found here may prove insightful for future modelling in the field.
@@ 5.3. Astrophysics dataset
Astrophysics constitutes an ideal testing environment for MvSR as it encompasses a diverse array of non-linear equations aimed at modeling observations of large-scale phenomenon within the universe. Classical SR has been widely used to characterise various relationships such as photometric redshift(Krone-Martins et al., 2014), the Hertzsprung–Russell diagram or the plane of elliptical galaxies (Graham et al., 2013). After this project was completed, we also became aware of (Tenachi et al., 2023), who approached a similar problem when dealing with the treatment of stellar streams.
@@
The largest catalogs of variability classifications contain millions of stellar objects (Watson et al., 2006; Eyer et al., 2023), providing researchers an access to numerous light curves of the same type of variability. One of the most studied types of extreme variability is supernovae (SNe), a catastrophic cosmic event caused by a star’s explosion. The physical processes in a supernova are highly complex, and we may not yet be able to directly model individual supernova light curves. This makes such astrophysical data an excellent case study for MvSR.
We choose three Supernova type Ia (SNIa) with a good time sampling from ZTF Data Release 17: SN2019fck, SN2018aye and SN2021mwb (Bellm et al., 2019; Malanchev et al., 2023). We use observations in $g$ and $r$ photometric filters independently, resulting in a dataset of six examples for MvSR. The light curve shape of SNIa in $g$ filter could be roughly described as an exponentially rising brightness followed by an exponential fading. Such behavior is typically well modeled by parametric equations found in the literature (Bazin et al., 2009; Villar et al., 2019; Sánchez-Sáez et al., 2021; Russeil et al., 2023). However, being observed in a redder $r$ filter, SNIa light curves have a secondary bump. This particular behaviour is much harder to describe with a simple parametric equation. We choose to use both $g$ and $r$ band to provide a wide variety of examples to MvSR.
We apply data quality cuts and select only data points with a signal to noise ratio greater than 20. The data was shifted so that the observed peak time is at 0 and the entire light curve was normalized by its maximum flux. Finally we consider only data points with a time ranging from -50 to 150 days from the observed peak.
@@
Noticeably, the second equation is already known as the  Bazin function (Bazin et al., 2009) and is widely used in the literature (Russeil et al., 2023)(Dai et al., 2018). However, the standard form includes a $t_{0}$ parameter appearing twice which is added to t, effectively encoding for a time shift. The form generated by MvSR is mathematically equivalent but uses parameters appearing only once. The first solution presented in Table 4 is characterized by an intricate exponential form. Despite lacking similar counterparts in the literature, it provides excellent fits and even outperforms the standard Bazin function (Figure 5). Moreover, since it uses only 3 parameters, additional ones such as $t_{0}$ or a scaling parameter could be added to make it more versatile. Finally the last equation in Table 4 requires 5 parameters and provides the most accurate description of the SNe. MvSR can produce solutions of arbitrary size with increasingly good fit, however we choose to limit it to 5 parameters maximum in order to prevent the overfitting.
@@
Overall MvSR was able to generate multiple good models to describe SNe behaviors. It recovered a solution from the literature and even proposed improved solutions in terms of goodness of fit and/or number of parameters used. However, the generated models struggle with the same problem as the equations from the literature: they don’t provide a simple description of the second bump of the SNIa in the $r$ band. It may also highlight a limitation of the current MvSR implementation, which doesn’t include parameter repetition, as stated in Section 2). Given that a repeated time shift parameter is standard in models used in the literature, this results highlights the importance of a more complete MvSR implementation for practical scientific applications.
@@ 6. Conclusions
We have tested this approach with four different challenging artificial benchmarks with added noise and composed of extreme situations where some of the parameters are set to 00 or the domain coverage is limited. When compared to the traditional approach we report that MvSR is capable of correctly retrieving the original expression in most configurations with a higher accuracy than SR, even in a presence of a strong noise. Additionally, we stress-tested our method by using real-world experiments from three different areas: chemistry, finance and astrophysics. MvSR was not only capable of recovering well known models from the literature but it also found new alternatives that are promising in these fields.
@@
In this context, a full implementation of MvSR would have much broader applications than the ones described here. We intend to further explore its ramifications in dedicated studies focused on the particularities of each science case. Nevertheless, in order to fully exploit its potential and popularize its use within a broad range of scientific areas, a user-friendly implementation of all the above mentioned functionalities is paramount.
@@ Acknowledgements
This project emerged from discussions during two SNAD workshops: SNAD-IV [^8] https://snad.space/2021/ and SNAD V [^9] https://snad.space/2022/, held in France, Briançon and Clermont Ferrand, respectively. ER, KM and EEOI thank the entire SNAD team for the insightful discussions and environment. ER acknowledges the ISITE grant from the ”AAP Séjour-recherche” program from Université Clermont-Auvergne (UCA). FOF acknowledges support from Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP), grant number 2021/12706-1 and CNPq through the grant 301596/2022-0. EEOI thanks Rafael S. de Souza and Alberto Krone-Martins for the introduction to Symbolic Regression. Finally we thank the pastéis for being a true source of inspiration.

### v3

@@ 5.2. Finance dataset
Financial markets exhibit complex emergent behaviors. We define the return as the difference between the price of an asset at time $t$ and time $t+1$ is $r_{t}=p_{t+1}-p_{t}$. At first approximation, we can describe the distribution of returns by a Brownian motion (Bachelier, 1900). The field of econophysics, which applies tools from statistical physics to study these stochastic processes, seeks to identify probability distribution functions that go beyond such Gaussian approximation. Those initial models led to the development of the renowned Black-Scholes equation (Black and Scholes, 1973), but neglected the significance of rare and extreme events due to their lack of fat tails. Indeed, the distribution of these specific events can be fitted with power-laws (Mantegna and Stanley, 1995).
@@
Modern models, as presented in (Mantegna and Stanley, 1999), focus on Lévy processes. These include, for example, the Gaussian distribution as well as the Cauchy distribution, the second being an already improved solution with its fat tails. Such models were proposed after studying the statistical properties that most data exhibit, such as the famous $S\&P500$ dataset, and then finding distributions that possess such properties. However, our Multiview approach enables us to identify a common distribution for all assets by considering each of them individually. In this section, we show that MvSR rediscover some presented distributions and propose new models that better fit the real return distributions.
@@
We analyze time series data consisting of multiple prices generated by various assets. Each asset’s price corresponds to the value of a share of a company on financial markets. We utilize a publicly available Kaggle dataset  [^3] https://www.kaggle.com/datasets/iveeaten3223times/massive-yahoo-finance-dataset containing data from $491$ companies. The first 10 companies are used as views for MvSR, while the remaining companies are only used for testing purposes. Values of assets are taken each day at open market time, over a period of 5 years starting from January $1^{st}$, $2018$. For each asset, we study the distribution of its returns with a sampling of $100$ bins of equal width. We normalize the data as described in Section 3.1. This type of data exhibits common statistical properties. A positive mean that corresponds to overall economic growth. A leptokurtic profile that indicates more extreme events than a normal distribution would produce, and a negative skewness due to an asymmetric shape that informs rare events are more likely to be crises than economic booms. The first and last properties are only observed for datasets covering long time periods (months or years) such as ours.
@@
Under those conditions, we explore multiple seeds, tree lengths (ranging from size 8 to 20) and operators (using various combinations of $exp$, $abs$ and $power$) to obtain multiple possible parametric solutions with MvSR. In Table 3 we present 6 parametric models generated by MvSR. We measure the performance of all the functions by fitting them to each asset, and display the median MSE value as a comparative metric in the last two columns of Table 3.
@@
Figure 4. Normalized (Section 3.1) distribution of returns for 2 assets, fitted by the Cauchy model and the best MvSR solution (Power-Laplace). The first asset is an example of a high MSE improvement in the usage of Power-Laplace compared to Cauchy.
Figure 4 shows the distributions of returns of some assets alongside the fit given by the best distributions, new and already found in the literature, that we obtain. We call the latter a Power-Laplace distribution  [^4] That we name after it’s functional composition $h(x)=g(f(x))$ with the Laplace distribution $g(x)=e^{-a\times|x|}$ and a power function $f(x)=x^{b}$.. It performs well because the power is always a root  [^5] We always find $10<b<$ 1., which accentuates the fat tail effect of the distribution, allowing a better fit in the tails while maintaining a good peaked fit in the center.
@@
Finally, we see that distributions found in the literature are relevant for the $S\&P500$, as they all perform much better than the Gaussian. However, apart from the Gaussian, their scores do not differ much. This shows that they are almost all equivalent fits for this aggregated index. On the other hand, their median scores for a variety of assets are much more disparate. In this second case, our algorithm produced much more efficient distributions ; distributions that outperformed all the others. This indicates that our multiview approach is an efficient strategy for finding a unique distribution that gives good fits for as many assets as possible. We illustrate this effect on Figure 4. The right panel shows how Cauchy and Power-Laplace both fit well the $S\&P500$ index. The left panel displays a special case selected for its high score difference between both models. This shows how a model like Cauchy fails to fit some distributions compared to the best model found by MvSR, which is more versatile.
We showed that applying MvSR to a set of assets all at once, rather than looking at them separately or using an aggregated index, is an efficient approach to gaining new insights into markets behavior and improving characterization of the stochastic process that governs them. The results presented here could be fine-tuned by adding more functions to the MvSR. For example, the Gamma function could allow us to generate more distributions such as the Gamma variance process (Madan et al., 1998) and Student’s t-distribution (Bouchaud and Potters, 2000), which are the general forms for the distributions we already found using MvSR. Finally, further investigation of the distributions found here may prove insightful for future modelling in the field.
@@ 5.3. Astrophysics dataset
Astrophysics constitutes an ideal testing environment for MvSR as it encompasses a diverse array of non-linear equations aimed at modeling observations of large-scale phenomenon within the universe. Classical SR has been widely used to characterize various relationships such as photometric redshift(Krone-Martins et al., 2014), the Hertzsprung–Russell diagram or the plane of elliptical galaxies (Graham et al., 2013). After this project was completed, we also became aware of (Tenachi et al., 2023), who approached a similar problem when dealing with the treatment of stellar streams.
@@
The largest catalogs of variability classifications contain millions of stellar objects (Watson et al., 2006; Eyer et al., 2023), providing researchers access to numerous light curves of the same type of variability. One of the most studied types of extreme variability is supernovae (SNe), a catastrophic cosmic event caused by a star’s explosion. The physical processes in a supernova are highly complex, and we may not yet be able to directly model individual supernova light curves. This makes such astrophysical data an excellent case study for MvSR.
We choose three Supernova type Ia (SNIa) with a good time sampling from ZTF Data Release 17: SN2019fck, SN2018aye and SN2021mwb (Bellm et al., 2019; Malanchev et al., 2023). We use observations in $g$ and $r$ photometric filters independently, resulting in a dataset of six examples for MvSR. The light curve shape of SNIa in $g$ filter could be roughly described as an exponentially rising brightness followed by an exponential fading. Such behavior is typically well modeled by parametric equations found in the literature (Bazin et al., 2009; Villar et al., 2019; Sánchez-Sáez et al., 2021; Russeil, E. et al., 2024). However, being observed in a redder $r$ filter, SNIa light curves have a secondary bump. This particular behavior is much harder to describe with a simple parametric equation. We choose to use both $g$ and $r$ band to provide a wide variety of examples to MvSR.
We apply data quality cuts and select only data points with a signal-to-noise ratio greater than 20. The data was shifted so that the observed peak time is at 0 and the entire light curve was normalized by its maximum flux. Finally, we consider only data points with a time ranging from -50 to 150 days from the observed peak.
@@
Noticeably, the second equation is already known as the  Bazin function (Bazin et al., 2009) and is widely used in the literature (Russeil, E. et al., 2024)(Dai et al., 2018). However, the standard form includes a $t_{0}$ parameter appearing twice which is added to t, effectively encoding for a time shift. The form generated by MvSR is mathematically equivalent but uses parameters appearing only once. The first solution presented in Table 4 is characterized by an intricate exponential form. Despite lacking similar counterparts in the literature, it provides excellent fits and even outperforms the standard Bazin function (Figure 5). Moreover, since it uses only 3 parameters, additional ones such as $t_{0}$ or a scaling parameter could be added to make it more versatile. Finally, the last equation in Table 4 requires 5 parameters and provides the most accurate description of the SNe. MvSR can produce solutions of arbitrary size with increasingly good fit, however we choose to limit it to 5 parameters maximum in order to prevent the overfitting.
@@
Overall, MvSR was able to generate multiple good models to describe SNe behaviors. It recovered a solution from the literature and even proposed improved solutions in terms of goodness of fit and/or number of parameters used. However, the generated models struggle with the same problem as the equations from the literature: they don’t provide a simple description of the second bump of the SNIa in the $r$ band. It may also highlight a limitation of the current MvSR implementation, which doesn’t include parameter repetition, as stated in Section 2). Given that a repeated time shift parameter is standard in models used in the literature, this results highlights the importance of a complete MvSR implementation for practical scientific applications.
@@ 6. Conclusions
We have tested this approach with four different challenging artificial benchmarks with added noise and composed of extreme situations where some parameters are set to 00 or the domain coverage is limited. When compared to the traditional approach, we report that MvSR is capable of correctly retrieving the original expression in most configurations with a higher accuracy than SR, even in a presence of a strong noise. Additionally, we stress-tested our method by using real-world experiments from three different areas: chemistry, finance and astrophysics. MvSR was not only capable of recovering well known models from the literature, but it also found new alternatives that are promising in these fields.
@@
In this context, a full implementation of MvSR would have much broader applications than the ones described here. We intend to further explore its ramifications in dedicated studies focused on the particularities of each science case. Nevertheless, in order to fully exploit its potential and popularize its use within a broad range of scientific areas, a user-friendly implementation of all the above-mentioned functionalities is paramount.
@@ Acknowledgements
This project emerged from discussions during two SNAD workshops: SNAD-IV [^7] https://snad.space/2021/ and SNAD V [^8] https://snad.space/2022/, held in France, Briançon and Clermont Ferrand, respectively. ER, KM and EEOI thank the entire SNAD team for the insightful discussions and environment. ER acknowledges the ISITE grant from the ”AAP Séjour-recherche” program from Université Clermont-Auvergne (UCA). FOF acknowledges support from Fundação de Amparo à Pesquisa do Estado de São Paulo (FAPESP), grant number 2021/12706-1 and CNPq through the grant 301596/2022-0. EEOI thanks Rafael S. de Souza and Alberto Krone-Martins for the introduction to Symbolic Regression. Finally we thank the Brazilian pastéis for being a true source of inspiration.